OpenStudy (anonymous):

PLEASE HELP Solve the rational equations: a) 4/x-2 - 3/x+1 = 8/x^2-x-2 b) 2x/x-2 - 4x-1/3x+2 = 17x+4/3x^2-4x-4 c) 5x+2/10x-3 = x-8/2x+3 d) x/3x-5 = 2/x-1 I've tried so many ways and have not been able to come up with an answer

OpenStudy (anonymous):

ill try, would you mind helping me on mine?

OpenStudy (er.mohd.amir):

take LCM on LHS and make factor of x2-x-2 cancel same factor

OpenStudy (er.mohd.amir):

LHS=4(x+1)-3(x-2) upon (x+1)*(x-2)

OpenStudy (anonymous):

Ok so lets see. Substitute x with 1 to begin with.

OpenStudy (anonymous):

that works btw, you can check the work if you want, it ends up being -2=-2

OpenStudy (er.mohd.amir):

x2-x-2 =x2-(2-1)x-2 =x2-2x+x-2 =x(x-2)+1(x-2) =(x-2)(x+1)

OpenStudy (er.mohd.amir):

cancel x-2 on both side